OpenStudy (anonymous):

What is the value of sin67.5? use half angle formula

OpenStudy (anonymous):

sin(67.5) = sin(135/2) Recall that sin²(x/2) = (1 - cos(x))/2 by the half angle formula. sin(x/2) = +/- √((1 - cos(x))/√2 In this case, sin(67.5) lies in Quad. 1 so sine is positive. sin(67.5) = √(1 - cos(135))/√2 = √(1 + √2/2)/√2 = √(√2/4 + 1/2) = 0.9239

OpenStudy (anonymous):

i was just missing the very last stepp thankk youu!